Find the 90% confidence interval for the population standard deviation given the following. n = 55, x = 11.49, s=2.34 and the distribution is normal.
Answer #1

We have given here,                                              
                                              
                                  
Sample standard deviation=2.34                                  
Sample size =55                                      
                                              
Degree of freedom =n-1=   54                                  
Level of significance given =1-0.90=0.1                                  
Chi square critical value for lower tail=38.1162                          
                                              
Chi square critical value for upper tail=72.1532                          
                                              
We can use here confidence interval formula for population standard deviation
(n 1)S2 2 χ(a/2 ) (n 1)S2 <σ< 2 Xi-(a/2)
|(55 1)(2.34)a (55 1) (2.34)2 38.1162 72.1532

=( 2.02, 2.79)   
                                              
                                              
                                              
                      
                                              
Lower confidence limit for population standard deviation=2.02                  
                                              
Upper confidence limit for Population standard deviation=2.79